# Phys4DGen: Physics-Compliant 4D Generation with Multi-Material Composition Perception

### v4

@@ 4.1.1 Implementation Details
Phys4DGen supports both a single image and a 3D model as input. Given a single image, we use LGM [37] to generate static 3D Gaussians. For material grouping, we render multiview images from the generated 3D Gaussians and apply SAM2 [29] to obtain cross-view consistent material mask maps. We used GPT-4o to identify the material for each material region in the image. For 4D dynamics generation, we perform physical simulation using MPM [34]. For each example, the simulation environment is configured based on the material information inferred by GPT-4o, and different external forces are applied according to the specific case, allowing the generation of physically plausible dynamic 4D sequences. All experiments were conducted on NVIDIA A40(48GB) GPU. For more detailed information on the experimental settings, please refer to the Appendix.
@@ 4.1.3 Baselines
Method PR $\uparrow$ OQ $\uparrow$ CLIP-T $\uparrow$ CLIP-T other $\uparrow$
DG4D [30] 5.9 6.25 0.98983 0.98536
STAG4D [45] 5.55 5.97 0.98813 0.98492
L4GM [31] 6.3 6.7 0.99242 0.99275
Ours 7.5 7.72 0.99459 0.99409
Table 1: Quantitative comparison in image-to-4D generation. PR evaluates the physical realism of the 4D content, OQ measures its overall quality, and CLIP and CLIP-T-other assesses its spatiotemporal consistency.
@@ 4.1.4 Metrics
Following previous works [49, 44], We use CLIP-T score which calculates the average cosine similarity between the CLIP embeddings of every two adjacent frames in rendered video from a given view. To further assess the spatiotemporal consistency, we render videos from the right, back, and left views to calculate the CLIP-T-other score. We also conduct a user preference study to evaluate the physical realism (PR) and overall quality (OQ) of the generated 4D content, with both metrics rated on a 10-point scale. More details on user study can be found in Appendix.
@@ 4.2 Showcase of 4D Generation Results
Fig. 3 visualizes the 4D content generated by Phys4DGen. For each example, we present the perceived material properties, including the density and Young’s modulus of both the object’s surface and internal regions. As shown, Phys4DGen effectively discerns the material compositions of 3D objects, such as differentiating the distinct physical attributes of carnations’ petals and stems, and recognizing the material differences in a doll’s fabric surface and polyester fiberfill interior. Furthermore, we render the corresponding 4D content from dynamically changing viewpoints. The results demonstrate Phys4DGen’s capability to generate physically realistic 4D content from a single image or 3D content. Please refer to the Appendix for additional visual results.
@@ 4.3 Comparison in Image-to-4D Generation
Fig. 4 presents a qualitative comparison of our Phys4DGen method with state-of-the-art (SOTA) image-to-4D generation methods. Generating 4D content using STAG4D and DG4D heavily depends on the quality of reference videos produced by video diffusion models. However, due to the inherently stochastic and data-driven nature of these models, it remains challenging to obtain reference videos that simultaneously adhere to physical laws and faithfully align the user’s intent. This limitation is evident in Fig. 4, where the 4D content generated by STAG4D and DG4D struggles with controllability and often violates fundamental physical principles. For example, in Fig. 4, the 4D content generated by DG4D [30] for a balloon tied to a wooden block moves upwards, violating gravity. In stark contrast, Fig. 4 clearly demonstrates that our generated 4D content achieves high fidelity and adheres physical laws, significantly outperforming baseline methods. This qualitative observation is further supported by the quantitative results presented in Tab. 1. As shown in the table, our method achieves the highest CLIP-T and CLIP-T-other scores, indicating that it is capable of generating spatiotemporally consistent 4D content. The user study results on physical realism (PR) and overall quality (OQ) indicate that participants found our generated 4D content more physically plausible and expressed a stronger preference for our results over the baselines. Additionally, our method offers superior controllability, enabling users to adjust external forces according to their specific needs—an aspect where baseline approaches fall short.
@@ 4.4 Comparison in 3D-to-4D Generation
Following PhysDreamer[48], we compare our results with real captured videos and simulations from other methods using space-time slices, where the vertical axis represents time and the horizontal axis shows a spatial slice of the object, as indicated by the red lines in the ”object” column. We use 90-frame video sequences to generate space-time slices. This representation enables an effective comparison of the magnitude and frequencies of the oscillatory motions generated by different methods. Fig. 5 demonstrates that the dynamics generated by our method more closely resembles that of real-world videos. As shown in Tab. 2, our method achieves the highest average scores in physical realism (PR), overall quality (OQ), and CLIP-T, significantly outperforming baseline methods.
@@
Table 2: Quantitative comparison in 3D-to-4D generation. Since PhysGaussians lacks a material property estimation stage, we exclude its runtime from our evaluation. Our method requires only 14.88s to material properties estimation.
These results validate the effectiveness of our proposed multi-material composition perception capability, which enables relatively accurate material property estimation within only 14.88 seconds, in contrast to the hour-level computation time required by baseline methods. This efficient perception contributes to the generation of 4D content that is not only physically plausible but also more preferred by human observers. More comparative results are provided in the Appendix.
@@ 4.5 Ablation Analysis
Material Grouping enables material editing by allowing different material parts to be independently controlled during simulation. As shown in Figure 7, we are able to simulate the melting effect specifically on the snowman’s body without affecting the rest of the object, such as the scarf. In contrast, without material grouping, the snowman is treated as a uniform material entity. As a result, the melting simulation affects the entire structure, including the scarf, which clearly violates physical plausibility. This demonstrates that multi-material perception is essential for fine-grained control and physically meaningful manipulation.
@@ 5 Conclusion
In this paper, we propose Phys4DGen, a physics-compliant 4D generation framework that effectively perceives complex multi-material compositions. By seamlessly integrating these perceptual capabilities with physical simulation, our approach enables intuitive and physically plausible 4D generation from a single image or a 3D input. To handle multi-material compositions, we propose the 3D Material Grouping module, which segments an object surface, represented by 3D Gaussians, into distinct material regions. Furthermore, the internal structure of the object is modeled through Physical Internal Structure Discovery. We distill extensive physical priors from GPT-4o to identify surface and internal material properties, which are then assigned to the 3D representation to construct a complete simulation object. Extensive experiments on synthetic and real-world datasets show that our approach generates physically realistic 4D content.

### v5

@@ 4.1.1 Implementation Details
Phys4DGen supports both a single image and a 3D model as input. Given a single image, we use LGM [43] to generate static 3D Gaussians. For material grouping, we render multiview images from the generated 3D Gaussians and apply SAM2 [35] to obtain cross-view consistent material mask maps. We used GPT-4o to identify the material for each material region in the image. For 4D dynamics generation, we perform physical simulation using MPM [40]. For each example, the simulation environment is configured based on the material information inferred by GPT-4o, and different external forces are applied according to the specific case, allowing the generation of physically plausible dynamic 4D sequences. All experiments were conducted on NVIDIA A40(48GB) GPU. For more detailed information on the experimental settings, please refer to Appendix A.1.
@@ 4.1.4 Metrics
Following previous works [56, 51], we use CLIP-T score which calculates the average cosine similarity between the CLIP embeddings of every two adjacent frames in rendered video from a given view. To further assess the spatiotemporal consistency, we render videos from the right, back, and left views to calculate the CLIP-T-other score. We also conduct a user preference study to evaluate the physical realism (PR) and overall quality (OQ) of the generated 4D content, with both metrics rated on a 10-point scale. More details on the user study can be found in Appendix A.3
@@ 4.2 Showcase of 4D Generation Results
Fig. 3 visualizes the 4D content generated by Phys4DGen. For each example, we present the perceived material properties, including the density and Young’s modulus of both the object’s surface and internal regions. As shown, Phys4DGen effectively discerns the material compositions of 3D objects, such as differentiating the distinct physical attributes of carnations’ petals and stems, and recognizing the material differences in a doll’s fabric surface and polyester fiberfill interior. Furthermore, we render the corresponding 4D content from dynamically changing viewpoints. The results demonstrate Phys4DGen’s capability to generate physically realistic 4D content from a single image or 3D content. Please refer to Appendix B.1 for additional visual results.
Method PR $\uparrow$ OQ $\uparrow$ CLIP-T $\uparrow$ CLIP-T other $\uparrow$
DG4D [36] 5.90 6.25 0.98983 0.98536
STAG4D [52] 5.55 5.97 0.98813 0.98492
L4GM [37] 6.30 6.70 0.99242 0.99275
Ours 7.50 7.72 0.99459 0.99409
Table 1: Quantitative comparison in image-to-4D generation. PR evaluates the physical realism of the 4D content, OQ measures its overall quality, and CLIP and CLIP-T-other assess its spatiotemporal consistency.
@@ 4.3 Comparison in Image-to-4D Generation
Fig. 4 presents a qualitative comparison of our Phys4DGen method with state-of-the-art (SOTA) image-to-4D generation methods. Generating 4D content using STAG4D and DG4D heavily depends on the quality of reference videos produced by video diffusion models. However, due to the inherently stochastic and data-driven nature of these models, it remains challenging to obtain reference videos that simultaneously adhere to physical laws and faithfully align with the user’s intent. This limitation is evident in Fig. 4, where the 4D content generated by STAG4D and DG4D struggles with controllability and often violates fundamental physical principles. For example, in Fig. 4, the 4D content generated by DG4D [36] for a balloon tied to a wooden block moves upwards, violating gravity. In stark contrast, Fig. 4 clearly demonstrates that our generated 4D content achieves high fidelity and adheres physical laws, significantly outperforming baseline methods. This qualitative observation is further supported by the quantitative results presented in Tab. 1. As shown in the table, our method achieves the highest CLIP-T and CLIP-T-other scores, indicating that it is capable of generating spatiotemporally consistent 4D content. The user study results on physical realism (PR) and overall quality (OQ) indicate that participants found our generated 4D content more physically plausible and expressed a stronger preference for our results over the baselines. Additionally, our method offers superior controllability, enabling users to adjust external forces according to their specific needs—an aspect where baseline approaches fall short.
@@ 4.4 Comparison in 3D-to-4D Generation
Following PhysDreamer[55], we compare our results with real captured videos and simulations from other methods using space-time slices, where the vertical axis represents time and the horizontal axis shows a spatial slice of the object, as indicated by the red lines in the ”object” column. We use 90-frame video sequences to generate space-time slices. This representation enables an effective comparison of the magnitude and frequencies of the oscillatory motions generated by different methods. Fig. 5 demonstrates that the dynamics generated by our method more closely resemble those of real-world videos.
@@
Table 2: Quantitative comparison in 3D-to-4D generation. Since PhysGaussians lacks a material property estimation stage, we exclude its runtime from our evaluation. Our method requires only 14.88s for property estimation.
As shown in Tab. 2, our method achieves the highest average scores in physical realism (PR), overall quality (OQ), and CLIP-T, significantly outperforming baseline methods. These results validate the effectiveness of our proposed multi-material composition perception capability, which enables relatively accurate material property estimation within only 14.88 seconds, in contrast to the hour-level computation time required by baseline methods. This efficient perception contributes to the generation of 4D content that is not only physically plausible but also more preferred by human observers. More comparative results are provided in Appendix B.2.
@@ 4.5 Ablation Analysis
Material Grouping enables material editing by allowing different material parts to be independently controlled during simulation. As shown in Figure 7, we are able to simulate the melting effect specifically on the snowman’s body without affecting the rest of the object, such as the scarf. In contrast, without material grouping, the snowman is treated as a uniform material entity. As a result, the melting simulation affects the entire structure, including the scarf, which clearly violates physical plausibility. This demonstrates that multi-material perception is essential for fine-grained control and physically meaningful manipulation. Additional experimental analyzes can be found in the Appendix B.4.
@@ 5 Conclusion
In this paper, we propose Phys4DGen, a physics-compliant 4D generation framework that effectively perceives complex multi-material compositions. By seamlessly integrating these perceptual capabilities with physical simulation, our approach enables intuitive and physically plausible 4D generation from a single image or a 3D input. To handle multi-material compositions, we propose the 3D Material Grouping module, which segments an object surface, represented by 3D Gaussians, into distinct material regions. Furthermore, the internal structure of the object is modeled through Physical Internal Structure Discovery. We distill extensive physical priors from GPT-4o to identify surface and internal material properties, which are then assigned to the 3D representation to construct a complete simulation object. Extensive experiments on synthetic and real-world datasets show that our approach generates physically realistic 4D content. Currently, our method is designed for single-object 4D generation. Extending it to handle multi-object scenarios is an exciting and challenging avenue for future work.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. 62276222.
Appendix A More Experimental Settings
A.1 More Implementation Details
Our method supports 4D generation from either a single image or an given 3D content. For single-image input, we first adopt LGM [43] to generate a static 3DGS representation from the image. For 3D content input, we directly utilize the provided model as the starting point. For the image-to-4D task, we render 29 images from different viewpoints around the generated 3D representation. For the 3D-to-4D task, we render 120 images from randomly sampled viewpoints based on the given 3D Gaussian representation. Based on the resulting image sequences, we use SAM2 [35] to extract material grouping masks with cross-view consistency. We set the occlusion threshold to 0.1. If the difference between the actual and estimated depth is less than this threshold, the Gaussian kernel is considered visible. For Gaussian kernels occluded from any viewpoint, we assign their material group to that of their nearest neighbor. For Gaussians that are occluded from all rendering viewpoints, we assign their material group based on the nearest visible Gaussian in spatial proximity. For internal physical structure discovery, we voxelize the object’s bounding boxes into a 32×32×32 grid. We use the GPT-4o to infer material properties and set the inference temperature to 0.8. After that, we extract the CLIP features of the segmented sub-images using CLIP ViT-B-16 [34] to perform CLIP Fusion. For all examples in the Physically-Compliant 4D Dynamics Generation, the material properties are initialized according to the material properties predicted by the previously described GPT-4o model. For the image-to-4D task, the foreground region is discretized into a $50^{3}$ Eulerian grid. We set 714 sub-steps between successive frames, corresponding to a duration of $5\times 10^{-5}$ seconds for each sub-step, allowing for the generation of 14 frames per second. For the 3D-to-4D task, the foreground region is discretized into a $64^{3}$ Eulerian grid. Between successive frames, 400 simulation sub-steps are performed, with the time step set consistent with the image-to-4D setting.
A.2 Prompting Details
We provide prompts for material reasoning, as shown in Fig. 8. By inputting these prompts along with the input image and its segmented sub-images into GPT-4o, we can infer the material type, density, Young’s modulus, Poisson’s ratio, and other physical properties.
Figure 8: Prompt used for material identification
Figure 9: Visualization results of more examples. From the input images, we generated the 4D content, including orange roses swaying in the wind, sandcastle collapsing, and a rubber duck being pressed.
A.3 User Study Details
To further evaluate the quality of the 4D content generated by our method, we conducted a user study involving 25 volunteer participants. During the experiment, each participant was asked to view a series of test examples, each generated by multiple existing methods including ours. To eliminate potential ordering bias, the 4D results from different methods were presented in a randomized sequence for each example. Participants independently rated each method on two criteria:
  • Physical Realism (PR): Whether the generated motion complies with real-world physical laws, such as whether the deformation aligns with material properties, whether the motion is stable, and whether any physically implausible behaviors occur.
  • Overall Quality (OQ): The overall perceptual quality of the animation, including visual coherence, structural consistency, and naturalness of material appearance.
Both metrics were rated on a 10-point scale (1 = lowest, 10 = highest). The evaluation was self-paced with no time constraints, and all participants completed the ratings independently to ensure unbiased subjective judgment.
Appendix B More Experimental Results
B.1 Visual Results
We provide visualization results for more examples, as shown in Fig. 9, including pink roses swaying in the wind, a sandcastle collapsing, a rubber duck being pressed and a basketball falling down under gravity. These results further highlight the generalization ability of our method, showcasing its capability to generate 4D content in diverse real-world scenarios.
Figure 10: More Comparison in 3D-to-4D Generation. We further tested various methods on the PhysDreamer dataset under large external forces, where our method exhibited dynamics most consistent with physical laws.
Figure 11: Failure Case. MLLMs misclassified both petals and leaves as plastic and set Young’s modulus to 1 GPa. This led to stiff 4D dynamics that do not match how real flowers bend under force.
B.2 More Comparisons in 3D-to-4D Generation
To further evaluate the physical realism of different physics-based methods under large external forces, we design three representative experimental scenarios using real-world objects: Carnations, Alocasia, and Telephone. We compare the performance of PhysGaussian, PhysDreamer, Physics3D, and our proposed method. As shown in Fig. 10, we present the corresponding temporal frame sequences along with space-time slices to analyze the oscillation frequency and dynamic response characteristics produced by each method. Experimental results show that PhysGaussian and PhysDreamer often generate overly distorted and unstable dynamics, lacking physical plausibility. Although Physics3D exhibits a certain degree of elastic recovery, its overall responsiveness remains limited. In contrast, our method demonstrates stronger structural preservation, more realistic material responses, and smoother temporal continuity across diverse scenarios. It is capable of producing physically consistent and stable 4D dynamics for objects with various shapes and material properties. These results further verify the generalization ability and robustness of our approach under complex external forces, offering a reliable solution for high-fidelity 4D content generation.
Figure 12: External Forces Analysis. 4D content is generated by applying different external forces $\mathbf{f}$.
B.3 Failure Case Analysis
Due to the inherent stochasticity of MLLMs, material properties estimation may occasionally be erroneous. As a failure case, we reference the Red Flower example in Fig. 4 of the paper. Due to current review constraints, this visualization cannot yet be shown, but it will be included in the Appendix. As shown in Fig. 11, the MLLM incorrectly classified both the petals and leaves as plastic, assigning them a Young’s modulus of 1 GPa. This resulted in a generated 4D dynamics with a rigid appearance, which clearly contradicts the expected soft and flexible behavior of real flowers under force. In contrast, the result shown in Fig. 4 demonstrates a physically consistent 4D dynamics generated based on a correct material prediction. We hypothesize this error was caused by visual ambiguity in the input image: the Red Flower appears artificial, leading the MLLM to misinterpret it as a plastic object and overestimate its stiffness.
B.4 Ablation Studies
The Ability for Fine-grained Controlling 4D Dynamics. Fig. 12 illustrates the 4D content generated under various external forces, showcasing the fine-grained controllability of Phys4DGen. In the first row, external force $\mathbf{f}_{1}=(1.0,0.0,0.0)$, directed to the right along the x-axis is applied, causing the orange rose to move to the right. In the second row, a larger force $\mathbf{f}_{2}=(2.0,0.0,0.0)$ is applied in the same direction, resulting in more intense motion. This demonstrates that Phys4DGen can control the strength of motion by adjusting external force. In the third row, a leftward force $\mathbf{f}_{3}=(-1.0,0.0,0.0)$, with the same magnitude as in the first row, is applied along the x-axis. As a result, the orange rose moves left under this external force, demonstrating that Phys4DGen can control the direction of motion To summarize, Phys4DGen allows fine-grained control of the dynamics in 4D content by adjusting external forces.
Hyperparameter Analysis for Internal Structure Discovery. During the internal structure discovery stage, we adopted the same set of heuristic-driven parameters across all scenarios, including: grid size = 32 (which controls particle filling density), confidence = 0.95 (used in the chi-squared test function to determine the distance threshold), and max particle number = 10 (the maximum number of particles allowed per Gaussian kernel).
Specifically, we conducted hyperparameter analysis over the following parameter combinations: confidence $\in\{0.75,0.85,0.95\}$, max particle number $\in\{10,50,100\}$ and grid size $\in\{16,32,50\}$. As shown in Fig. 13, we primarily investigated the influence of confidence and max particle number on the simulation results. The experiments show that increasing either parameter leads to more frequent bounces of the basketball after impact, indicating enhanced elasticity. We attribute this behavior to the increased number of surface particles introduced by higher parameter values. Since the predicted Young’s modulus of the surface material is higher than that of the internal material, the elevated surface particle ratio results in an overall increase in elastic behavior. Additionally, in the Fig. 14, we performed an ablation study on grid size. The results show that increasing the grid size leads to higher toughness during simulation. This observation aligns with the behavior of the MPM framework, where increased particle density typically contributes to greater material stiffness.
Figure 13: Different Settings of Confidence and Max Particles Parameters. Higher parameters cause more bounces after landing, implying greater elasticity and stiffness.
Figure 14: Different Settings of Grid Size Parameters. Larger grid sizes cause more simulated toughness, matching MPM trend where higher particle density raises material stiffness.
[…]
Top-1 Material/Ratio Snow/100% Fabric/74% Felt/100%
Top-1 $E$ ($\log_{10}(Pa)$)/Ratio 6/48% 7.699/24% 7.699/68%
Top-1 $\rho$ ($kg/m^{3}$)/Ratio 100/94% 1500/64% 200/78%
Top-1 $\nu$/Ratio 0.3/82% 0.4/68% 0.3/74%
$E$ Mean/Median 5.738/5.699 8.219/7.699 7.917/7.699
$\rho$ Mean/Median 113/100 1369/1500 346/200
$\nu$ Mean/Median 0.287/0.3 0.374/0.4 0.307/0.3
Table 3: Statistics of 50 MLLMs estimations on the Snowman example. Reported metrics include Top-1 predicted material, Young’s modulus ($E$, in $\log_{10}(Pa)$), density ($\rho$, in $kg/m^{3}$), and Poisson’s ratio ($\nu$), along with mean and median values across multiple runs.
[…]
Carnation (petal/stem/pot) 6.00 / 7.176 / 9.301 6.00 / 7.301 / 9.176 6.00 / 7.00 / 9.176
Basketball 7.00 7.00 7.00
Table 4: Comparison of Young’s modulus ($E$, in $\log_{10}(Pa)$) predicted by different MLLMs on the same objects.
B.5 Reliability Analysis of MLLMs
We conducted an experiment on the Snowman example (Fig. 7), performing 50 independent material property estimation, and reported the statistical results in Tab. 3. The results show that the material predictions from the MLLM do not follow a uniform random distribution, but instead exhibit strong preference patterns—indicating that the model tends to predict certain material properties more frequently and confidently. We also compared the material indentification across different MLLMs for the same object, and summarized the predicted Young’s modulus values for each surface material in Tab. 4. The results show a high degree of consistency across models. This inter-model agreement suggests that MLLMs may have already internalized shared knowledge about material properties during pretraining, reflecting real-world material semantics to some extent. As demonstrated in all experiments in this paper, the 4D content generated based on MLLM-estimated material properties exhibits plausible visual dynamics and strong physical consistency. These results collectively indicate that MLLMs can serve as effective material property estimators, offering reasonably accurate predictions that support downstream physical simulation tasks.
Appendix C Preliminaries
C.1 3D Gaussian Splatting
Unlike NeRF  [28, 5, 29], which uses implicit representations, 3D Gaussian Splatting (3DGS) employs explicit representations through a set of anisotropic Gaussian kernels [15], which can be interpreted as particles in space, enables physical simulations to simulate the particles’ motion. The explicit representations also enable superior reconstruction quality and faster rendering speed. Specifically, 3DGS represents the 3D scene using a set of anisotropic Gaussian kernels defined by center position (mean) $\mathbf{x}\in\mathbb{R}^{3}$, covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{7}$, opacity $\alpha\in\mathbb{R}$, and color $\mathbf{c}\in\mathbb{R}^{3}$. 3D Gaussian kernel’s spatial distribution can be expressed by:
$G(\mathbf{x})=e^{-{\cfrac{1}{2}}(\mathbf{x})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x})}.$ (7)
During the rendering process, the 3D Gaussians will be projected onto the image plane as 2D Gaussians. Sequentially, the pixel color $\mathbf{C}(\mathbf{\mathbf{r}})$ rendered from ray $\mathbf{r}$ can be computed by point-based volume rendering technique:
$\mathbf{C}(\mathbf{r})=\sum_{i\in\mathcal{N}}\mathbf{c}_{i}\sigma_{i}\prod_{j=1}^{i}(1-\sigma_{j}),$ (8)
where $\sigma_{i}=\alpha_{i}G(\mathbf{x}_{i})$, and $\mathcal{N}$ denotes the number of Gaussian kernel along ray $\mathbf{r}$. Because of this explicit representation, tracking the state of the Gaussian kernels becomes straightforward.
C.2 Material Point Method
Continuum mechanics studies the deformation and motion behavior of materials under forces. Motion is typically represented by the deformation map $\mathbf{x}=\phi(\mathbf{X},t)$, which maps from the undeformed material space $\omega^{0}$ to the deformed world space $\omega^{t}$. The deformation gradient $\mathbf{F}=\frac{\partial\phi}{\partial\mathbf{X}}(\mathbf{X},t)$ describes how the material deforms locally. MPM is a simulation method that combines Lagrangian particles with Eulerian grids and has demonstrated its ability to simulate various materials. In MPM, each particle $p$ carries various physical properties $\boldsymbol{\theta}_{p}^{t}$ at time step $t$, including mass $m_{p}$, density $\rho_{p}$, volume $V_{p}$, Young’s modulus $E_{p}$, Poisson’s ratio $\nu_{p}$, velocity $\mathbf{v}_{p}^{t}$, deformation gradient $\mathbf{F}_{p}^{t}$ and velocity gradient $\mathbf{C}_{p}^{t}$. The grid $i$ is used for computing intermediate results. MPM operates within a loop that includes particle-to-grid (P2G) transfer, grid operations, and grid-to-particle (G2P) transfer. In the particle-to-grid (P2G) stage, MPM transfers momentum and mass from particles to grids:
$(m\mathbf{v})_{i}^{t+1}$ $=\sum_{p}w_{ip}\left[m_{p}\mathbf{v}_{p}^{t}+m_{p}\mathbf{C}_{p}^{t}(\mathbf{x}_{i}-\mathbf{x}_{p}^{t})\right],$ (9)
$m_{i}^{t+1}$ $=\sum_{p}w_{ip}m_{p},$ (10)
where $w_{ip}$ is the B-spline kernel that measures the distance between particle $p$ and grid $i$. After P2G stage, we perform grid operations:
$\mathbf{v}_{i}^{t}$ $=(m\mathbf{v}_{i})^{t}/{m_{i}^{t}}$ (11)
$\mathbf{f}_{i,in}^{t}$ $=-\sum_{p}\mathbf{P}(\mathbf{F}_{p})(\mathbf{F}_{p})^{T}\nabla w_{ip}\mathbf{v}_{p}^{0}$ (12)
$\mathbf{v}_{i}^{t+1}$ $=\mathbf{v}_{i}^{t}+\Delta t\left(\mathbf{f}_{i,in}+\mathbf{f}_{ex}\right)/m_{i}$ (13)
$\mathbf{v}_{i}^{t+1}$ $=BC(\mathbf{v}_{i}^{t+1})$ (14)
where BC refers to boundary condition. Then we transfer the results back to particles in the grid-to-particle (G2P) stage:
$\mathbf{v}_{p}^{t+1}$ $=\sum_{i}w_{ip}\mathbf{v}_{i}^{n+1},$ (15)
$\mathbf{x}_{p}^{t+1}$ $=\mathbf{x}_{p}^{t}+\Delta t\mathbf{v}_{p}^{t+1}.$ (16)
The velocity $\mathbf{v}_{p}^{t+1}$ and position $\mathbf{x}_{p}^{t+1}$ are updated using semi-implicit Euler method. Then, we update velocity gradient $\mathbf{C}_{p}^{t+1}$ and deformation gradient $\mathbf{F}_{p}^{t+1}$:
$\mathbf{C}_{p}^{t+1}$ $=\frac{4}{\Delta\mathbf{x}^{2}}\sum_{i}w_{ip}\mathbf{v}_{i}^{t+1}(\mathbf{x}_{i}-\mathbf{x}_{p}^{t})^{T},$ (17)
$\mathbf{F}_{p}^{t+1}$ $=\left(\mathbf{I}+\Delta t\mathbf{C}_{p}^{t+1}\right)\mathbf{F}_{p}^{t}.$ (18)
By following these three stages, we complete a simulation step.
C.3 External Forces.
In physical simulations, external forces—such as gravity—directly influence the motion and deformation of a continuum system. In this paper, we apply two types of external forces. The first type directly modifies the particle’s velocity, such as setting the velocity of all particles to simulate the translation of the continuum. The second type indirectly affects the particle’s velocity by applying forces $\mathbf{f}$. For example, gravity can simulate the continuum’s fall. Given a force $\mathbf{f}$, we apply Newton’s second law and time integration to compute the particle’s velocity at the next time step:
$\mathbf{a}_{p}^{t}=\frac{\mathbf{f}}{m_{p}^{t}},\quad\mathbf{v}_{p}^{t+1}=\mathbf{v}_{p}^{t}+\mathbf{a}_{p}^{t}\Delta t,$ (19)
where $\mathbf{a}_{p}^{t}$ denotes the acceleration of particle $p$ at time step $t$, and $\Delta t$ denotes the time interval between time step $t$ and $t+1$. By adjusting the external forces, we can control the motion and deformation of the object, thereby controlling the dynamics of the 4D content. Additionally, since material grouping for 3D objects has been implemented in Sec. 3.2, we can precisely apply external forces to any material part of the object without relying on manual region selection, achieving user-friendly interaction and fine-grained control.